OpenStudy (anonymous):

a company offers a starting yearly salary of $37,000 with raises of $3,500 per year. Find the total salary over a ten year period

OpenStudy (anonymous):

229500 I think

OpenStudy (anonymous):

37,000 + 3500(10) = total salary for ten years 37000 + 35000 = 72,000 you might want to get a second opinion on this :/

OpenStudy (anonymous):

Well, its asking for the total salary, which I take to mean the total amount of money he made. So the first year would be 37000 + 3500, since he gets a raise. That is to be added to the second year, where he would make 37000 + 3500, + 3500, since he had two raises. And so on. He didn't make 72000 over ten years or he, his family and his cats would be long dead. But who knows.

OpenStudy (anonymous):

would it be 37,000(10) + 3500(10) = total salary ?

OpenStudy (anonymous):

@Luis_Rivera .....help.....please

OpenStudy (anonymous):

@jim_thompson5910 .....help

jimthompson5910 (jim_thompson5910):

He made 37,000 the first year He made 37,000+3,500 = 40,500 the second year He made 37,000+2*3,500 = 44,000 the third year etc etc He made 37,000+9*3,500 = 68,500 the tenth year

jimthompson5910 (jim_thompson5910):

So your goal is to add up the numbers 37000 40500 44000 .... ... 68500

OpenStudy (anonymous):

it says find the total salary for ten years

jimthompson5910 (jim_thompson5910):

A short cut is to use this formula Sn = n*(a1 + an)/2

jimthompson5910 (jim_thompson5910):

Sn = n*(a1 + an)/2 S10 = 10*(a1 + a10)/2 S10 = 10*(37000+68500)/2 S10 = 10*(105500)/2 S10 = 10*52750 S10 = 527500

jimthompson5910 (jim_thompson5910):

So the total salary is $527,500

OpenStudy (anonymous):

oh...I see now.....jim, your good...thanks :)

jimthompson5910 (jim_thompson5910):

thx for the compliment lol and yw

OpenStudy (anonymous):

You are good. Yeah I think thats right!